OpenStudy (anonymous):

What is the graph of the function? y = 2 · 2x

OpenStudy (anonymous):

is it y=2-2x ?

OpenStudy (anonymous):

if that's so. Find where it intersects x and y -axis and it will give you the graph

OpenStudy (anonymous):

No let me put the options :p

OpenStudy (anonymous):

so it is multiplied? y=2(2x)

OpenStudy (anonymous):

These are the options it gives me

OpenStudy (anonymous):

can you draw how is the function exactly?

OpenStudy (anonymous):

y = 2 · 2x this is all it gave me

OpenStudy (anonymous):

what that point in the midle means?

OpenStudy (anonymous):

multiplication

OpenStudy (anonymous):

so it's y=4x

OpenStudy (anonymous):

So which graph would it be?

OpenStudy (anonymous):

then none of the graph represent's it, because it's a line